We have replaced both optics and the patch lead. Light levels have improved and we are no longer seeing errors increment on the link.
Posted Jul 23, 2019 - 23:02 BST
Scheduled maintenance is currently in progress. We will provide updates as necessary.
Posted Jul 23, 2019 - 22:01 BST
Our monitoring system has detected a potential issue with one of our datacentre uplink fibres in London. We will be carrying out emergency maintenance to replace/clean optical fibre modules and the fibre patch lead starting at 22:00 BST this evening.
This maintenance will affect customers with Ethernet services delivered over BT Wholesale or SSE circuits in Telehouse North and the expected downtime is up to 10 minutes whilst components are replaced. As this maintenance is being carried out on a core network device all other customers that transit this device will be 'At Risk' during the work. Where possible we will be rerouting traffic away from the device.
Posted Jul 23, 2019 - 16:02 BST
This scheduled maintenance affected: Core (Core Network Devices and Routing) and Connectivity (Ethernet Services (Fibre, EFM, GEA)).